OpenStudy (anonymous):

The vertical asymptote of the function f(x) = 3 log(x + 3) is x = what

OpenStudy (solomonzelman):

you shifted the function \(\large\color{slate}{ \displaystyle y=\log(x) }\) which way ? (disregard the coefficient in front for this problem)

OpenStudy (anonymous):

idk I added 3? @SolomonZelman

OpenStudy (solomonzelman):

yes, \(\large\color{slate}{ \displaystyle f(x+c) }\) is a shift to the left by c units from \(\large\color{slate}{ \displaystyle f(x) }\).

OpenStudy (anonymous):

so x is 3? @SolomonZelman

OpenStudy (solomonzelman):

where would the asymptote for \(\normalsize \color{blue}{ \displaystyle y=\log(x) }\) normally be?

OpenStudy (solomonzelman):

take some number that approaches zero from the right side (such that it is getting 0.1 to 0.023 to 0.0007 to 0.00001 to 0.0000...0001 ) doesn't have to use those values particularly, but the point is that it is getting closer and closer to zero.

OpenStudy (solomonzelman):

would it make sense to think of such a thing when the number is closer and close approaching zero, as of \(\large\color{black}{ \displaystyle \frac{1}{\infty } }\) ?

OpenStudy (solomonzelman):

and infinity symbol is there to represent a very large number. because 0.0000...0001 is essentially `1/(really big number)`

OpenStudy (solomonzelman):

am I making sense

OpenStudy (solomonzelman):

?

OpenStudy (solomonzelman):

(so far)

OpenStudy (anonymous):

yeah most of it does @SolomonZelman

OpenStudy (solomonzelman):

yes, now you know the rule for logarithms that \(\large\color{black}{ \displaystyle \log(a^b)=b\cdot \log(a) }\) and you know that \(\large\color{black}{ \displaystyle \frac{1}{n}=n^{-1} }\)

OpenStudy (solomonzelman):

So, as you are taking the log of very small number, watch what happens...

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \log({\rm very~~small~~number}) }\) \(\large\color{black}{ \displaystyle \log(\frac{1}{{\rm very~~big~~number}}) }\) \(\large\color{black}{ \displaystyle \log(~~({\rm very~~big~~number})^{-1}~~) }\) \(\large\color{black}{ \displaystyle (-1)\cdot \log({\rm very~~big~~number}) }\) \(\large\color{blue}{ \displaystyle -\log({\rm very~~big~~number}) }\)

OpenStudy (solomonzelman):

as this "small number" getting smaller and smaller (saying as it closer and closer approaches zero), the big number (at this time) which is on the bottom is getting larger and larger. so in the result, as this "small number" is closer to zero, the "very big number" is going more and more towards infinity. AS WE APPLY THE RULES, we have -log( very big number) meaning, we get something along the lines of \(\large\color{black}{ \displaystyle -\log(\infty) }\) WHICH would be just as good as saying \(\large\color{black}{ \displaystyle -\infty }\)

OpenStudy (solomonzelman):

now, we can watch on the graph, https://www.desmos.com/calculator/txcoxfbguq

OpenStudy (solomonzelman):

you can see it never hits zero, but the close x to zero - the y is going more towards \(\large\color{black}{ \displaystyle -\infty }\)

OpenStudy (solomonzelman):

when you shift this function y=log(x) 3 units left, you are shifting it's asymptote of x=0 and your new asymptote is?